OpenStudy (anonymous):

23. Beth scored 21 points, 24 points, 15 points, and 21 points in the first four games of the basketball season. How many points must she score in her next game to average 20 points per game for the first 5 games? (4 points)

OpenStudy (anonymous):

(21 + 24 + 15 + 21 + X ) = 100 (in order to get 20 avg, she needs exactly 100 points over 5 games) X = 100 - 21 - 15 - 24 - 21

OpenStudy (anonymous):

X=19 then?